can someone solve this word problem thank you
An actor invests some money at 5%, and $32000 more than twice the amount the amt at 9%. The total annual interest earned from the investment is 24960. How much did he invest at each amount?use the six step method

Answer by mananth(16946)

An actor invests some money at 5%,--------------x
amount at 9% --------= 2x+32000
5%x+(32000+2x)*9%=24960
multiply by 100
5x+(32000+2x)*9 = 2496000
5x+18x+288000=2496000
23x= 2496000-288000
23x=2208000
x=96000-----------------at5%
96000+32000= 128000-----------at 9%